# Efficiency-Effectiveness Reranking FLOPs for LLM-based Rerankers

### v1

@@ 5 Experimental Results and Analysis
Q1: What are the LLM-based reranker performances under the new metrics?
Table 2 reports the efficiency–effectiveness trade-offs of a broad set of LLM-based rerankers under the proposed RPP and QPP metrics. Overall, these systems perform poorly once computation is taken into account. Across all LLMs and methods, TourRank with Llama-3.1-8B-Instruct secures the highest NDCG on both DL19 and DL20 at the cost of almost the lowest RPP and QPP. Beyond this result, several additional insights emerge:
Pointwise methods dominate the RPP and QPP metrics across different LLMs and datasets. pointwise.yes_no of Flan-T5-large yields the highest RPP of 72.67 (DL19) and 68.3 (DL20) and achieves the maximum QPP of 111 queries/PetoFLOPs. The variant pointwise.qlm obtains the similar QPP metric but 5 10 worse RPP points. These methods obtain 10% to 30% relative NDCG gains over the baseline BM25 with negligible FLOPs consumption compared with other LLM-based rerankers.
@@
Pairwise and listwise methods are intensely FLOP-hungry. Allpair sorting, though it delivers the highest NDCG on Flan-T5-xl (0.713), issues 9,900 LLM calls per query; its RPP collapses to around 0.10, and it processes barely 0.15 queries per peta FLOPs, making large-scale deployment impractical. Heapsort and bubblesort based variants cut the number of calls by roughly 90%, yet remain about orders of magnitude less efficient than pointwise methods on both RPP and QPP.
Q2: Do the estimated FLOPs reflect the measured FLOPs?
Figure 1 shows the relationship between the estimated and real FLOPs on DL19 for models of various sizes. The comparison contains both decoder-only and encoder-decoder architectures, providing a comprehensive view of scaling trends. In both cases, the estimated and real FLOP counts scale with the model size, reflecting the expected rise in computational cost with increasing model parameters. The linear pattern across models illustrates that the estimated FLOPs correlate linearly with the real FLOPs and is consistent across model families and architectural types, suggesting that our FLOPs estimator is accurate and reliable. The close alignment between the two quantities provides empirical justification for the FLOPs estimator described in Section 3, affirming its reliability as a proxy when real measurements are unavailable.
Q3: How does latency relate to the FLOP counts?
Figure 2 shows the relationship between latency and FLOP counts for two representative models: Qwen-7B (a decoder-only architecture) and Flan-T5-XXL (an encoder-decoder architecture), evaluated on the DL19 dataset. For both models, we observe that latency increases in accordance with the number of FLOPs. Importantly, the estimated FLOP counts exhibit a correlation with latency that closely mirrors the relationship between real FLOP counts and latency. This alignment indicates that our FLOPs estimator approximates computational cost accurately and can serve as a reliable predictor of real-world latency trends. This means that the estimator can be used to anticipate inference time without requiring direct hardware profiling, which is particularly useful when comparing models in a platform-agnostic setting or during early-stage architecture design.
Q4: What is the relationship between prompt length and FLOPs?
Figure 3 shows the relationship between prompt length and FLOPs. As expected, both the estimated and actual FLOPs increase with longer prompts, reflecting the greater computational cost required to process more input tokens. Notably, the estimated FLOPs show a strong correlation with prompt length that closely mirrors the pattern observed in the actual FLOPs. As the prompt becomes longer, the estimator reliably tracks the resulting increase in computation, consistent with what is observed empirically. This result provides additional validation for the robustness of our FLOPs estimator, demonstrating its ability to respond to changes in input length in the same way real FLOPs do.
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
@@
Figure 1: Linear trends between estimated and real FLOPs for decoder (left) and encoder-decoder (right) models of various sizes on DL19. The same is observed for the DL20 dataset.
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
@@
Figure 2: Latency in milliseconds increases with FLOPs on Qwen-7B (left) and Flan-T5-XXL (right)
(a) Decoder (Qwen)
(b) Encoder-Decoder (T5)
@@ 6 Conclusion and Future Works
Due to limitation of existing metrics in evaluating the efficiency-effectiveness tradeoff of large language models as rerankers, we propose two metrics RPP and QPP to quantify the model performance. In addition, we provide a calculator through a closed-form and interpretable formula to compute the FLOPs and validate this estimation through experiments on existing decoder-only and encoder-decoder model architectures. The estimated FLOP count exhibits a strong linear correlation with the actual measured values, allowing it to approximate real-world computational cost even without model execution. Future work includes conducting a linear regression between the real and estimated FLOP counts to refine our estimation and adapting to more advanced architectures.
@@
While our proposed FLOP-normalized metrics and estimator provide a hardware-agnostic and interpretable approach to evaluating the efficiency–effectiveness tradeoff of LLM-based rerankers, there are some limitations. The FLOP estimation relies on model architecture specifications and assumes consistent implementation across different frameworks, which may not hold in practice due to library-level optimizations or kernel differences. Although the estimator shows strong linear correlation with actual FLOP measurements, the approximation may be less accurate for models with more advanced architectures such as mixture-of-experts. While FLOPs offer a more stable proxy than latency or token counts, they do not capture other real-world constraints such as memory bandwidth, energy consumption, or inference-time variability under dynamic system loads.

### v2

@@ 5 Experimental Results and Analysis
Q1: Can E2R-FLOPs overcome the limitations of existing efficiency proxies?
Existing efficiency proxies for LLM-based rerankers, such as latency Jin et al. (2025), number of LLM calls (i.e.,, forward passes) Zhuang et al. (2024), and token counts Chen et al. (2025b), are weak surrogates for actual compute. Latency is confounded by hardware, parallelism, etc.; the same algorithm can appear faster or slower across different platforms. LLM-call counts discard model size and sequence length, for instance, one call on a 70B model is orders of magnitude more expensive than one call on a 3B model, yet both count as “1.” Token counts ignore model size and the prefill-decode asymmetry (quadratic attention during prefill versus near‑linear growth during decoding), so equal token totals can yield very different FLOPs.
[…]
FLOPs, in contrast, is an intrinsic, hardware‑agnostic measure of work. Normalizing reranking metrics and throughput by FLOPs (RPP/QPP) enables fair, interpretable comparisons across different LLMs. Additionally, it aligns with trends in measured FLOPs and latency (Section 3, Figure 1, Figure 2).
Q2: What are the LLM-based reranker performances under the E2R-FLOPs?
Table 2 reports the efficiency–effectiveness trade-offs of a broad set of LLM-based rerankers under the proposed RPP and QPP metrics. Overall, these systems perform poorly once computation is taken into account. Across all LLMs and methods, TourRank with Llama-3.1-8B-Instruct achieves the highest NDCG on both DL19 and DL20, albeit at the cost of almost the lowest RPP and QPP. Beyond this result, several additional insights emerge.
[…]
@@
Pairwise and listwise methods are intensely FLOP-hungry. Allpair sorting, although it delivers the highest NDCG on Flan-T5-xl (0.713), issues 9,900 LLM calls per query; its RPP collapses to around 0.10, and it processes barely 0.15 queries per petaFLOPs, making large-scale deployment impractical. Heapsort- and bubblesort-based variants cut the number of calls by roughly 90%, yet remain about orders of magnitude less efficient than pointwise methods on both RPP and QPP.
(a) Decoder
(b) Encoder-Decoder
@@
Figure 1: Linear trends between estimated and measured FLOPs for decoder (left) and encoder-decoder (right) models of various sizes on DL19. The same is observed for the DL20 dataset.
Q3: Do the estimated FLOPs reflect the measured FLOPs?
Figure 1 shows the relationship between the estimated and measured FLOPs on DL19 for models of various sizes. The comparison contains both decoder-only and encoder-decoder architectures, providing a comprehensive view of scaling trends. In both cases, the estimated and measured FLOP counts scale with the model size, reflecting the expected rise in computational cost with increasing model parameters. The linear pattern across models illustrates that the estimated FLOPs correlate linearly with the measured FLOPs and are consistent across model families and architectural types, suggesting that our FLOPs estimator is accurate and reliable. The close alignment between the two quantities provides empirical justification for the FLOPs estimator described in Section 3, affirming its reliability as a proxy when real measurements are unavailable.
Q4: How does latency relate to the FLOP counts?
Figure 2 shows the relationship between latency and FLOP counts for two representative models: Qwen-7B (a decoder-only architecture) and Flan-T5-XXL (an encoder-decoder architecture), evaluated on the DL19 dataset. For both models, we observe that latency increases in accordance with the number of FLOPs. Importantly, the estimated FLOP counts exhibit a correlation with latency that closely mirrors the relationship between measured FLOP counts and latency. The Pearson correlation coefficients between latency and estimated FLOP counts are 0.88 for Qwen-7B and 0.94 for Flan-T5-XXL. This alignment indicates that our FLOPs estimator approximates computational cost accurately and can serve as a reliable predictor of real-world latency trends. This means that the estimator can be used to anticipate inference time without requiring direct hardware profiling, which is particularly useful when comparing models in a platform-agnostic setting or during early-stage architecture design.
Q5: What is the relationship between prompt length and FLOPs?
Figure 3 shows the relationship between prompt length and FLOPs. As expected, both the estimated and actual FLOPs increase with longer prompts, reflecting the greater computational cost required to process more input tokens. Notably, the estimated FLOPs exhibit a strong correlation with prompt length, which closely mirrors the pattern observed in the actual FLOPs. The Pearson correlation coefficient between prompt length and FLOP counts is 1. As the prompt becomes longer, the estimator reliably tracks the resulting increase in computation, consistent with what is observed empirically. This result provides additional validation for the robustness of our FLOPs estimator, demonstrating its ability to respond to changes in input length in a manner consistent with measured FLOPs.
(a) Decoder
(b) Encoder-Decoder
@@
Figure 2: Latency in milliseconds increases with FLOPs on Qwen-7B (left) and Flan-T5-XXL (right). The Pearson correlation coefficients between latency and estimated FLOP counts are 0.88 for Qwen-7B and 0.94 for Flan-T5-XXL.
(a) Decoder
(b) Encoder-Decoder
@@ 6 Conclusion and Future Works
Due to the limitations of existing metrics in evaluating the efficiency-effectiveness tradeoff of large language models as rerankers, we propose two metrics, RPP and QPP, to quantify the model performance. In addition, we provide a calculator based on a closed-form and interpretable formula to compute the FLOPs, and validate this estimation through experiments on existing decoder-only and encoder-decoder model architectures. The estimated FLOP count exhibits a strong linear correlation with the actual measured values, allowing it to approximate real-world computational cost even without model execution. Future work includes conducting a linear regression between the measured and estimated FLOP counts to refine our estimation and adapting to more advanced architectures.
@@
The FLOP estimation relies on model architecture specifications and assumes consistent implementation across different frameworks, which may not hold in practice due to library-level optimizations or kernel differences. Although the estimator shows strong linear correlation with actual FLOP measurements, the approximation may be less accurate for models with more advanced architectures in the future. While FLOPs offer a more stable proxy than latency or token counts, they do not capture other real-world constraints such as memory bandwidth, energy consumption, or inference-time variability under dynamic system loads.